# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v2

@@ D.3 Proof of Theorem 3
$\pi^{\mathrm{proj}}:=T(\pi^{\mathrm{warm}})=-A_{\mathrm{warm}}^{-1}G_{\mathrm{warm}}.$
Let $\widehat{A}$ and $\widehat{G}$ be the stage-2 (BPTT/Monte Carlo) estimators of $A_{\mathrm{warm}}$ and $G_{\mathrm{warm}}$ used to form the deployable stage-2 control
$\widehat{\pi}_{\mathrm{agg,mix}}:=-\widehat{A}^{-1}\widehat{G}$.
Let $\pi^{\star}_{\mathrm{blind}}$ denote a locally optimal interior deployable $\theta$-blind policy, so in particular it satisfies the fixed-point (stationarity) identity
$\pi^{\star}_{\mathrm{blind}}=T(\pi^{\star}_{\mathrm{blind}}).$
Step 1 (Estimator-to-population projection stability). Assume we are in the stable inversion regime (as enforced in practice by ridge/diagnostics) so that the hypotheses of Proposition 1 apply on $D$ with $A=A_{\mathrm{warm}}$ and $G=G_{\mathrm{warm}}$. Then there exists a constant $C_{2}>0$ (depending only on the uniform inverse bound $\kappa$ and uniform $L^{\infty}$ bounds on the relevant $G$ terms) such that
$\|\widehat{\pi}_{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}},$ (97)
where one convenient choice is
@@
Step 2 (Slab-wise contraction for the population projection map). Assume $\pi^{\mathrm{warm}},\pi^{\star}_{\mathrm{blind}}\in U$ and let $\rho^{\ast}<1$ be the slab contraction constant of Proposition 2. Since $\pi^{\mathrm{proj}}=T(\pi^{\mathrm{warm}})$ and $\pi^{\star}_{\mathrm{blind}}=T(\pi^{\star}_{\mathrm{blind}})$, for each slab $D_{k}$ we have
$\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{k}=\|T(\pi^{\mathrm{warm}})-T(\pi^{\star}_{\mathrm{blind}})\|_{k}\leq\rho(\tau_{k})\,\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\rho^{\ast}\,\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}.$ (98)
@@
$r^{\mathrm{warm}}_{\mathrm{FOC,mix}}:=A_{\mathrm{warm}}\pi^{\mathrm{warm}}+G_{\mathrm{warm}},\hskip 18.49988pt\varepsilon^{\mathrm{mix}}_{\mathrm{warm}}:=\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{L^{2}(\mu)},\hskip 18.49988pt\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}:=\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{k}.$
@@
$\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}=\pi^{\mathrm{warm}}-(-A_{\mathrm{warm}}^{-1}G_{\mathrm{warm}})=A_{\mathrm{warm}}^{-1}\bigl(A_{\mathrm{warm}}\pi^{\mathrm{warm}}+G_{\mathrm{warm}}\bigr)=A_{\mathrm{warm}}^{-1}r^{\mathrm{warm}}_{\mathrm{FOC,mix}}.$
@@
$\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}\leq\|A_{\mathrm{warm}}^{-1}\|_{L^{\infty}(D)}\,\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{k}\leq\kappa\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}.$ (99)
@@
$\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}+\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\kappa\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}+\rho^{\ast}\,\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}.$
Since $\rho^{\ast}<1$, we can close slab-wise:
$\|\pi^{\mathrm{warm}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\frac{\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}.$ (100)
@@
$\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{k}\leq\frac{\rho^{\ast}\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm},k}.$ (101)
Step 5 (Global bound and completion). Sum over slabs using $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k}\|f\|_{k}^{2}$:
$\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{L^{2}(\mu)}\leq\frac{\rho^{\ast}\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm}}.$
@@
$\|\widehat{\pi}_{\mathrm{agg,mix}}-\pi^{\star}_{\mathrm{blind}}\|_{L^{2}(\mu)}\leq\|\widehat{\pi}_{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\|_{L^{2}(\mu)}+\|\pi^{\mathrm{proj}}-\pi^{\star}_{\mathrm{blind}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}}+\frac{\rho^{\ast}\kappa}{1-\rho^{\ast}}\,\varepsilon^{\mathrm{mix}}_{\mathrm{warm}},$
which is the claimed slab-wise residual-based $\theta$-blind policy-gap bound. ∎
Appendix E Implementation details for Section 3
This appendix provides reproducible step-by-step templates for the methods in Section 3. The high-level pipeline is summarized in Figure LABEL:fig:sec3-pipeline-sideways. For compactness we present one template per subsection of Section 3.
E.1 Stage 1 (PG–DPO) template for Section 3.1
@@ Appendix E Implementation details for Section 3
    Sample latent environment parameter. Sample $\theta\sim q$ inside the simulator (unseen by $\pi_{\varphi}$). (Variant: sample $\theta^{(i)}\sim q$ independently per episode; both are unbiased for $\nabla_{\varphi}J(\varphi)$.)
@@
    Simulate Euler rollouts. For each episode $i$, simulate the Euler scheme in (45)–(46) under the $\theta$-blind policy $\pi_{\varphi}$ and collect terminal utilities $\{U(X_{T}^{(i)})\}_{i=1}^{M}$.
@@
    Backpropagation through time (BPTT). Compute the Monte Carlo gradient estimator
@@
E.2 Stage 2 (P–PGDPO projection; mixed-moment $q$-aggregation) template for Section 3.2
Stage 2 is a post-processing map: given a warm-up $\theta$-blind policy $\pi_{\varphi^{\mathrm{warm}}}$, it estimates Pontryagin sensitivity objects by Monte Carlo and constructs a deployable projected control on a working-domain sample $z\sim\mu$.
@@
Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; working-domain sampler $\mu$ on $D$; budgets $(M_{z},M_{\theta},M_{\mathrm{MC}})$.
@@
    For each query state $z_{j}$, sample latent parameters. Sample $\{\theta_{\ell}\}_{\ell=1}^{M_{\theta}}\sim q$.
@@
    For each frozen $\theta_{\ell}$, estimate costates at $z_{j}$. For each $\ell=1,\dots,M_{\theta}$:
@@
    Solve the projection (mixed-moment aggregation). Whenever $\widehat{A}_{t}(z_{j})$ is invertible and the solve is numerically stable, compute the deployable projected control
@@
    which matches (59).
E.3 Coupling I: residual/control-variate projection (Section F.1)
@@
Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; query state(s) $z=(t,x,y)\sim\mu$; and stage 2 projection ingredients $(\widehat{A}_{t}(z),\widehat{G}_{t}^{\mathrm{mix}}(z))$ constructed as in Section E.2.
@@
    $\widehat{\pi}^{\mathrm{agg,mix}}(z)\;\leftarrow\;\pi_{\varphi^{\mathrm{warm}}}(z)-\big(\widehat{A}_{t}(z)\big)^{-1}\widehat{r}_{\mathrm{FOC}}(z).$
E.4 Coupling II: interactive distillation (Section F.2)
This subsection records an implementation template for interactive distillation: the projected output from stage 2 is used as a teacher signal during stage 1 training via the mixed objective (106). The teacher is built from the mixed-moment projected rule (possibly evaluated in residual form for variance reduction).
@@
Student parameters $\varphi$; teacher refresh interval $K$; distillation schedule $\lambda(n)$; working-domain sampler $\mu$.
@@
      Stage 1 update (PG–DPO step). Perform one PG–DPO update step on $J(\varphi)$ as in Section E.1.
@@
      Teacher refresh (every $K$ steps). If $n\bmod K=0$:
@@
        Set $\varphi^{-}\leftarrow\varphi$ (lagged copy).
@@
        For each $z_{j}$, run stage 2 under $\pi_{\varphi^{-}}$ (mixed-moment aggregation) to compute a projected teacher $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(z_{j})$. (In practice we compute it in residual form around $\pi_{\varphi^{-}}$ as in (105).)
@@
        Optionally filter states using diagnostics (Section E.5) and update the buffer:
@@
        Apply a gradient step to minimize the proximity term $\|\pi_{\varphi}(z)-\mathrm{stopgrad}(\pi^{\mathrm{teach}})\|^{2}$ with coefficient $\lambda(n)$, consistent with (106).
E.5 Engineering notes and stabilizers
This subsection collects practical stabilizers that we found helpful for reliable training and projection in high dimensions.
  • Antithetic sampling for $\theta$. When $q$ is symmetric (e.g. Gaussian in a latent normal parameterization), sample $\theta$ in antithetic pairs by drawing $z\sim\mathcal{N}(0,I)$ and using (z,z)(z,-z) to construct $(\theta^{+},\theta^{-})$. This reduces the variance of $q$-averaged quantities and typically improves the stability of stage 2 diagnostics on the working domain.
  • Blockwise Monte Carlo and robust aggregation. To control rare-tail domination, split Monte Carlo replications into $B$ blocks and compute blockwise averages of costate-driven ingredients (e.g. $\widehat{A}_{t}^{\theta}(z)$ and $\widehat{G}_{t}^{\theta}(z)$). Aggregate across blocks using a robust statistic such as the median or median-of-means, which makes the projection less sensitive to outlier trajectories.
  • Curvature/denominator stability checks. Because the projection map $(A,G)\mapsto-A^{-1}G$ can be sensitive to near-singularity of $A$, monitor the conditioning of $\widehat{A}_{t}$ (or failure rates of the linear solve). When diagnostics indicate ill-conditioning, skip projection-guided updates at that state or increase Monte Carlo budgets locally.
  • Residual magnitude as a reliability diagnostic. For the residual form, compute $\widehat{r}_{\mathrm{FOC}}(z)=\widehat{A}_{t}(z)\pi_{\varphi^{\mathrm{warm}}}(z)+\widehat{G}_{t}^{\mathrm{mix}}(z)$. Small $\|\widehat{r}_{\mathrm{FOC}}(z)\|$ indicates approximate satisfaction of the mixed-moment aggregated first-order condition at $z$ and empirically correlates with more reliable teacher targets.
  • Diagnostics-based teacher selection on the working domain. Rather than applying distillation on all sampled $\{z_{j}\}\sim\mu$, keep only states that pass a reliability predicate. In practice, filter using residual-magnitude thresholds together with stable linear-solve diagnostics to prevent a small subset of pathological states from contaminating the teacher buffer.
  • $\lambda$ schedule and safeguards. Use a warm-up period with $\lambda=0$ (pure PG–DPO) and increase $\lambda$ only after stage 2 diagnostics on the working domain are stable. To prevent the teacher term from dominating the ex–ante objective, cap the effective coefficient via

### v3

@@ E.3 Proof of Theorem 3
$\pi^{\mathrm{proj}}:=\mathcal{T}(\pi^{\mathrm{warm}})=-A_{\mathrm{warm}}^{-1}g_{\mathrm{warm}}.$
Let $\widehat{A}$ and $\widehat{g}$ be the stage-2 (BPTT/Monte Carlo) estimators of $A_{\mathrm{warm}}$ and $g_{\mathrm{warm}}$ used to form the deployable projected control
$\widehat{\pi}^{\mathrm{agg,mix}}:=-\widehat{A}^{-1}\widehat{g}$,
where $\widehat{A}$ is understood as the effective matrix after any inversion stabilizers (ridge/diagnostics/skip) so that $\widehat{A}$ is well-defined and invertible on the working domain. Let $\pi^{\star,\mathrm{blind}}$ denote a locally optimal interior deployable $\theta$-blind policy, so in particular it satisfies the fixed-point (stationarity) identity
$\pi^{\star,\mathrm{blind}}=\mathcal{T}(\pi^{\star,\mathrm{blind}}).$
Step 1 (Estimator-to-population projection stability). To apply Proposition 1 with $A=A_{\mathrm{warm}}$ and $g=g_{\mathrm{warm}}$, we assume we are in the stable inversion regime enforced by the stage-2 stabilizers, i.e.
$\|\widehat{A}-A_{\mathrm{warm}}\|_{L^{\infty}(\mathcal{D})}\leq(2\kappa)^{-1},$ (102)
where $\kappa$ is such that $\|A_{\mathrm{warm}}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq\kappa$. Define the $L^{2}$ estimation error level
@@
Then Proposition 1 yields $\|\widehat{A}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq 2\kappa$ and
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\|_{L^{2}(\mu)}\leq 2\kappa\,\|\widehat{g}-g_{\mathrm{warm}}\|_{L^{2}(\mu)}+2\kappa^{2}\,\|g_{\mathrm{warm}}\|_{L^{\infty}(\mathcal{D})}\,\|\widehat{A}-A_{\mathrm{warm}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}},$ (103)
where one may take, for instance,
$C_{2}:=2\kappa+2\kappa^{2}\,\|g_{\mathrm{warm}}\|_{L^{\infty}(\mathcal{D})}.$
@@
Step 2 (Slab-wise contraction for the population projection map). Assume $\pi^{\mathrm{warm}},\pi^{\star,\mathrm{blind}}\in\mathcal{U}$ and let $\varrho_{*}<1$ be the slab contraction constant of Proposition 2. Since $\pi^{\mathrm{proj}}=\mathcal{T}(\pi^{\mathrm{warm}})$ and $\pi^{\star,\mathrm{blind}}=\mathcal{T}(\pi^{\star,\mathrm{blind}})$, for each slab $\mathcal{D}_{k}$ we have
$\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{k}=\|\mathcal{T}(\pi^{\mathrm{warm}})-\mathcal{T}(\pi^{\star,\mathrm{blind}})\|_{k}\leq\varrho(\tau_{k})\,\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}\leq\varrho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}.$ (104)
@@
$r_{\mathrm{FOC,mix}}^{\mathrm{warm}}:=A_{\mathrm{warm}}\pi^{\mathrm{warm}}+g_{\mathrm{warm}},\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{L^{2}(\mu)},\hskip 18.49988pt\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{k}.$
@@
$\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}=A_{\mathrm{warm}}^{-1}r_{\mathrm{FOC,mix}}^{\mathrm{warm}}.$
@@
$\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}\leq\|A_{\mathrm{warm}}^{-1}\|_{L^{\infty}(\mathcal{D})}\,\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{k}\leq\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (105)
@@
$\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}\leq\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}+\varrho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}.$
Since $\varrho_{*}<1$, we can close slab-wise:
$\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}\leq\frac{\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (106)
@@
$\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{k}\leq\frac{\varrho_{*}\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$ (107)
Step 5 (Global bound and completion). Sum over slabs using $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k=1}^{L}\|f\|_{k}^{2}$:
$\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{L^{2}(\mu)}\leq\frac{\varrho_{*}\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}.$
@@
$\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\|_{L^{2}(\mu)}\leq\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\|_{L^{2}(\mu)}+\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{L^{2}(\mu)}\leq C_{2}\,\delta_{\mathrm{BPTT}}+\frac{\varrho_{*}\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}},$
which is the claimed slab-wise residual-based $\theta$-blind policy-gap bound (under the stable inversion condition (102)). ∎
Appendix F Implementation details for Section 3
This appendix provides reproducible step-by-step templates for the methods in Section 3. The high-level pipeline is summarized in Figure 4. For compactness we present one template per subsection of Section 3.
Conventions. Throughout, the deployed policy is $\theta$-blind. The latent parameter $\theta\sim q$ is sampled only inside the simulator. Stage 2 computes projection ingredients on a working domain $\mathcal{D}\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m}$ with reference measure $\mu$.
F.1 Stage 1 (DPO) template for Section 3.1
@@ Appendix F Implementation details for Section 3
    Sample latent parameter(s) inside the simulator. Sample $\theta\sim q$ and keep it frozen along the rollout(s). (Variant: sample $\theta^{(i)}\sim q$ independently per episode; both are unbiased for the discretized gradient of $J(\varphi)$.)
@@
    Simulate Euler rollouts. For each episode $i$, simulate the Euler scheme for (39)–(40) under the $\theta$-blind policy $\pi_{\varphi}$, and compute terminal utilities $\{U(X_{T}^{(i)})\}_{i=1}^{M}$.
@@
    BPTT (pathwise gradient). Compute the Monte Carlo gradient estimator
@@
F.2 Stage 2 (Pontryagin projection; mixed-moment $q$-aggregation) template for Section 3.2
Stage 2 is a post-processing map: given a warm-up $\theta$-blind policy $\pi_{\varphi^{\mathrm{warm}}}$, it estimates Pontryagin sensitivity objects by Monte Carlo and constructs a deployable projected control on queried working-domain states $z\sim\mu$.
@@
Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; working-domain sampler $\mu$ on $\mathcal{D}$; budgets $(M_{z},M_{\theta},M_{\mathrm{MC}})$.
@@
    For each query state $z_{j}$, sample frozen parameters. Sample $\{\theta_{\ell}\}_{\ell=1}^{M_{\theta}}\sim q$.
@@
    For each frozen $\theta_{\ell}$, estimate costate blocks at $z_{j}$. For each $\ell=1,\dots,M_{\theta}$:
@@
    Solve the projection (deployable $\theta$-blind control). Whenever $\widehat{A}_{t}(z_{j})$ is numerically invertible and passes stability diagnostics, compute
@@
    which matches (53). Otherwise, fall back to $\pi_{\varphi^{\mathrm{warm}}}(z_{j})$ (or a ridge-stabilized solve; see Appendix F.5).
F.3 Coupling I: residual/control-variate projection (Section G.1)
@@
Warm-up policy $\pi_{\varphi^{\mathrm{warm}}}$; query state(s) $z=(t,x,y)\sim\mu$; and stage 2 ingredients $(\widehat{A}_{t}(z),\widehat{g}_{t}^{\mathrm{mix}}(z))$ from Appendix F.2.
@@
    $\widehat{\pi}^{\mathrm{agg,mix}}(z)\;\leftarrow\;\pi_{\varphi^{\mathrm{warm}}}(z)-\big(\widehat{A}_{t}(z)\big)^{-1}\widehat{r}_{\mathrm{FOC}}(z),$
    with ridge/diagnostics/skip as needed (Appendix F.5).
F.4 Coupling II: interactive distillation (Section G.2)
Interactive distillation amortizes the stage 2 projected control into the policy network. A lagged policy copy $\pi_{\varphi^{-}}$ is used to construct a teacher on a slower timescale, and the student $\pi_{\varphi}$ is trained against a mixed objective (112).
@@
Student parameters $\varphi$; teacher refresh interval $n_{\mathrm{refresh}}$; distillation schedule $\lambda(n)$; working-domain sampler $\mu$.
@@
      Stage 1 update (DPO step). Perform one DPO ascent step on $J(\varphi)$ as in Appendix F.1.
@@
      Teacher refresh (every $n_{\mathrm{refresh}}$ steps). If $n\bmod n_{\mathrm{refresh}}=0$:
@@
        Set $\varphi^{-}\leftarrow\varphi$.
@@
        For each $z_{j}$, run stage 2 under $\pi_{\varphi^{-}}$ (mixed-moment aggregation) to compute the projected teacher $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(z_{j})$ (typically in residual form (111)).
@@
        Optionally filter unreliable states using diagnostics (Appendix F.5) and update the buffer:
@@
        Apply one optimizer step for the teacher proximity term in (112) (treating $\pi^{\mathrm{teach}}$ as $\mathrm{stopgrad}$-fixed).
F.5 Engineering notes and stabilizers
This subsection collects practical stabilizers that are useful for reliable training and projection.
  • Antithetic sampling for $\theta$. When $q$ is symmetric (e.g. Gaussian in a latent-normal parameterization), sample $\theta$ in antithetic pairs. This reduces the variance of $q$-averaged quantities and improves stage 2 concentration.
  • Blockwise Monte Carlo and robust aggregation. Split Monte Carlo replications into blocks and aggregate blockwise estimates using robust statistics (median / median-of-means) to reduce tail domination.
  • Curvature/denominator checks for $\widehat{A}_{t}$. Monitor conditioning of $\widehat{A}_{t}(z)$ (e.g. eigenvalue floors / condition numbers). If unstable, skip projection at $z$ or use ridge:
    $\widehat{A}_{t,\lambda}(z):=\widehat{A}_{t}(z)+\lambda I_{d},\hskip 18.49988pt\widehat{\pi}_{\lambda}(z):=-\widehat{A}_{t,\lambda}(z)^{-1}\widehat{g}_{t}^{\mathrm{mix}}(z).$
  • Residual magnitude as a reliability diagnostic. Track $\widehat{r}_{\mathrm{FOC}}(z)=\widehat{A}_{t}(z)\pi^{\mathrm{warm}}(z)+\widehat{g}_{t}^{\mathrm{mix}}(z)$. Small $\|\widehat{r}_{\mathrm{FOC}}(z)\|$ empirically correlates with reliable teacher targets.
  • Diagnostics-based teacher selection. For distillation, keep only $z$ that pass stability predicates (invertibility/conditioning + moderate residual), to prevent noisy stage 2 targets from contaminating the teacher buffer.
  • Scheduling $\lambda$ (teacher loss). Use an initial warm-up with $\lambda=0$ (pure DPO), then increase $\lambda$ only after stage 2 diagnostics stabilize. Optionally cap the effective coefficient via